You have used Cepheid variable to measure the distance to a nearby spiral galaxy. The spiral galaxy is 5.0 Mpc (megaparsecs) away. Three years ago, a Type 1a supernova was observed in this galaxy, and had a peak flux of 6.2 × 10-14W m-2.


You have just discovered a new Type 1a supernova in a very distant galaxy. This new supernova stayed bright for the same time as the one in the nearby spiral galaxy. The flux of the new supernova was 1.3 × 10-18W m-2.


What is the distance to the new supernova (in Mpc)?:


1
Expert's answer
2021-07-20T09:45:57-0400

If we don't take the cosmologic expansion into account, we assume the flux to be proportional to "r^{-2}" , so "\dfrac{I_1}{I_2} = \dfrac{r_2^2}{r_1^2}, \; r_2 = 218r_1 = 1.1\;\mathrm{Gpc}."


But if we take the expansion into account, we should write "I = \dfrac{L}{4\pi r_2^2(1+z)^2}" . For such distances the redshift is approximately 0.2, if we calculate it in form of "z = Hr_2\/c" , therefore, "(1+z)^2 \approx 1.5," "(1+z)\approx 1.24" and we should correct the estimate of "r_2" from the by "1\/1.24" and get the value of 880 Mpc.